OpenStudy (anonymous):

HELP! MEDAL! Which set of three numbers could represent the lengths of the sides of a right triangle? A. 7, 24, 25 B. 8, 9, 10 C. 9, 11, 14 D. 15, 18, 21

OpenStudy (anonymous):

can u please put a link if there is a picture

OpenStudy (anonymous):

What do we know about a right triangle we know that they have a 90 degree angles right and with a right triangle they form into a right triangle by a line of 180 degree

OpenStudy (anonymous):

yash

OpenStudy (anonymous):

There is no link for this picture......that's why I need help. I don't understand.

OpenStudy (anonymous):

is there a pic, if so why not copy and past the pic?

OpenStudy (anonymous):

paste*

OpenStudy (anonymous):

There isn't a picture at all.

OpenStudy (anonymous):

he said there is none

OpenStudy (anonymous):

Ahh then it is a word problem okay Do you know a formula for a right triangle?

OpenStudy (anonymous):

I think I'm supposed to use the Pythagorean Theorem.

OpenStudy (anonymous):

it is c my friend :D:D:DD

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

:P yw

OpenStudy (anonymous):

do i get a medal?????| dw:1426103892118:dw|

OpenStudy (anonymous):

Are you sure? I think you're right.

OpenStudy (anonymous):

@Destinyrules1234 it is against the code of conduct to give out Direct answers, it is best to hear from me than ambassador of openstudy or the Moderator

OpenStudy (anonymous):

yas i aam

OpenStudy (anonymous):

WAT IT IS AHHHHHHHHHHHHHHHHHHHHHHHH

OpenStudy (anonymous):

wait why dont i get medal D:\[D:\]

OpenStudy (anonymous):

I was working it out while she said it but okay. I wont do it in the future! Thank you! It's right.......according to my work.

OpenStudy (anonymous):

IM A BOY!

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

OMG LEMME GET A MEDA: plz

OpenStudy (anonymous):

YAYYYYYYYYYYYYYYYYYYYYYYYYY

OpenStudy (anonymous):

lol........okay. Yes I know that formula and @Destinyrules1234 I thought I gave u one

OpenStudy (anonymous):

you just fill in the multiple choices into the eqaution

OpenStudy (anonymous):

so we know that A is the short leg of the triangle

OpenStudy (anonymous):

now u did

OpenStudy (anonymous):

b is the long leg which is the left side of triangle

OpenStudy (anonymous):

if u want anymore help @Sexy_einstein just tag meeeeeeeeeeeeee

OpenStudy (anonymous):

THANK YOU @Destinyrules1234

OpenStudy (anonymous):

and @mitchal

OpenStudy (anonymous):

and the longer leg is C which is the hypotenuse

OpenStudy (anonymous):

np @Sexy_einstein :D

OpenStudy (anonymous):

youve been sooooo much help to me. Im so confusing

OpenStudy (anonymous):

bad new is C is incorrect answer cause 9^2+11^2= 14^2 you do the problem and it became 202=196 it does not equal to each other which therefore is incorrect

OpenStudy (anonymous):

lets do D shall we so 15^2+ 18^2= 21^2 225+324= 441 549=441

OpenStudy (anonymous):

therefore its incorrect

OpenStudy (anonymous):

lets do b

OpenStudy (anonymous):

8^2+9^2=10^2 64+81=100 145=100 is incorrect so A is the correct answers cause 7 times 7 is 49 so 49 + 576 cause 24 times 24 is 576 = to 25 times 25 is 625 49 Plus 576 is 625 and it is equal to 625 so A is correct

OpenStudy (anonymous):

@Sexy_einstein

OpenStudy (anonymous):

destiny_rule is incorrect It is A